### v1

@@ 1 Introduction
Symmetric Uncertainty has been introduced in [CUP:Press_et_al:1988], and it is defined as
@@
where $X,Y$ are categorical random variables, $H(\cdot)$ is the entropy, and $H(\cdot,\cdot)$ is the joint entropy.
Generalization of SU for the multivariate case, as well as its properties, and, in particular, its potential to define a distance metric among categorical random variables have been explored in [Entropy:Gomez-2021, ref-Gomez2023, IS:Sosa-2018, SBMAC:Sosa-2018]. Indeed, [ref-Gomez2023] bring us to realize that entropic correlation is an example of similarity.
@@ 2 Background: Entropic Correlation and Symmetric Uncertainty
Let us consider a fixed probability space $(D,\Omega_{D},P)$, and let $X:D\to\Sigma_{X}$ be a categorical (non-numeric, nominal or cualitative) random variable with possible values $\{x_{1},\ldots,x_{k}\}$ and probability mass function $P_{X}$. The entropy $H$ of the variable $X$ is a measure of the uncertainty in predicting the value of $X$, or size of the uncertainty on $X$, and is defined as [Renyi-1961]:
$H(X):=-\sum_{i}P_{X}(x_{i})\log_{2}(P_{X}(x_{i})),$ (2)
where $P_{X}(x_{i})$ is the prior probability of the value $x_{i}$ of $X$. Any base could be used for the log, but base 2 is used for historical reasons connected to information theory [Shannon-1948]. $H(X)$ can also be interpreted as a measure of the amount of information a discrete random variable $X$ requires for describing its behavior; or more compactly, as the variety inherent to $X$.
Given another categorical random variable $Y$, the conditional entropy $H(X|Y)$ quantifies the amount of information needed to describe the outcome of $X$ given that the value of $Y$ is known. It is defined as follows:
$H(X|Y):=-\sum_{j}\left[P_{Y}(y_{j})\sum_{i}P(x_{i}|y_{j})\log_{2}(P(x_{i}|y_{j}))\right],$ (3)
where $P(x_{i}|y_{j})$ is the posterior probability of a value $x_{i}$ for the variable $X$ given that the value of variable $Y$ is $y_{j}$.
@@
For more details, we refer the reader to [FIMU:Bouda], page 15.
Mutual Information. Using entropy and conditional entropy introduced above, the mutual information  [^1] In certain literature mutual information is denoted as $I(X;Y)$; see [ref-Fullwood2023]. $MI(X|Y)$ is introduced [MKP:Quinlan:1993]. Mutual Information measures the reduction in uncertainty about the value of $X$ when the value of $Y$ is known, and is defined as:
@@
and since $H(X,Y)$ is symmetric by direct substitution of $X,Y$ into equation (1), it results that $MI(X|Y)$ is symmetric as well, a quite convenient property for a paired measure. A formal proof can be found, for example, as Theorem 1 in [ref-Fullwood2023].
Symmetric Uncertainty. On the other hand, like with other informational association measures, MI tends to increase its value when the number of categories in $X$ and/or $Y$ increases, that is, it is biased towards high cardinality features [ML:Quinlan-1986]. Therefore, seeking to compensate for such bias, MI has to be normalized by the sum of entropies of the features. This measure, called Symmetric Uncertainty [CUP:Press_et_al:1988] is expressed as:
@@
As a final observation regarding $SU$, we recall from [IS:Sosa-2018] the following results.
@@
$H({\cal Q}|{\cal R}):=-\sum_{Q\in{\cal Q}}\sum_{R\in\cal R}P(Q\cap R)\log\left(\frac{P(Q\cap R)}{P(R)}\right).$
@@
Lemma 2.2 (Lemma 9.1.5 in [Book:Viana-Oliveira]).
@@ 3.1 Measuring similarity on a dataset: between observations and between variables
A set of data, or dataset, consists of $m$ observations taken across $n$ variables. Each observation is displayed as a row of $n$ individual values. Each set of values at the same position (column) constitute measurements of a variable, see Table 1.
@@
Looking row-wise at the dataset, any two observations (rows) can be compared to determine their similarity. If two rows are equal throughout all their values, we say that they are in full similarity or equivalently, we say that the distance between these observations is zero. If not all the values on the two rows are equal, several techniques are available to compute distances in this $n$-dimensional space, where a mix of data types is not unusual; see for instance Suárez-Díaz et al [UG:Suarez-2020].
We can also look at the same dataset in a column-wise manner [ref-Gomez2023]. In exploratory experiments and in model construction, one tries to establish how the value in one column (the outcome or class) results from the values in other columns, revealing a concomitance or correlation between sets of values.
@@ 20 R I L H D N
Table 2 presents data example for the personality traits of students applying for an internship. The $SU$ values between each variable and the gotHired class are:
  • SU(Neatness, GotHired) = 0.0535
  • SU(Creativity, GotHired) = 0.4627 (largest computed SU)
  • SU(Punctuality, GotHired) = 0.0535
  • SU(IQuotient, GotHired) = 0.0535
  • SU(AttentionType, GotHired) = 0.0192
@@ 3.2 Entropic equivalence and the metric structure
Less standardized is the concept of a similarity metric. Here we recall the definition given by Chen et al in [TCS:Chen-2009].
@@
In [TCS:Chen-2009], the authors also gave conditions for a similarity metric to yield a distance metric. More concretely, they established a precise relationship between normalized distances and similarity metrics. Here, normalized means that both the distance and the similarity map into $[0,1]$. For the reader convenience, let us recall their result here.
Lemma 3.2 (Corollary 1 in [TCS:Chen-2009]).
@@
Let $\cal C$ be a set of categorical random variables defined on a sample space $D$. If $X\colon D\to\Sigma_{X}$ and $Y\colon D\to\Sigma_{Y}$ are two elements in $\cal C$}, we say that they are indiscernible if they have the same histogram, modulo a permutation of the categories.
Let us recall that, given $X\colon D\to\Sigma_{X}$, the histogram is just the graph of the probability mass function $\mu_{X}\colon\Sigma_{X}\to\mathbb{R}$ given by
$P_{X}(S)=P(X^{-1}(S))=\sum_{x\in S}P(X=x).$
Thus, in the previous definition, we are asking for the graphs of $P_{X}$ and $P_{Y}$ to coincide after, if necessary, relabeling or reordering the categories in one of the sets $\Sigma_{X}$ or $\Sigma_{Y}$. In particular, for $X$ and $Y$ to be indiscernible, their codomains $\Sigma_{X}$ and $\Sigma_{Y}$ must have the same cardinality. In the following, we will refer to the indiscernibility condition as the existence of a bijection $\Sigma_{X}\to\Sigma_{Y}$ and that $P_{X}=P_{Y}$.
The indiscernibility just defined introduces an equivalence relation in $\cal C$. Let us denote by $\overline{{\cal C}}$ the space of equivalence classes. Notice that any two equivalence classes may have the same number of categories or a different number of categories.
@@
Example of indiscernibles. Consider variable $X_{1}$ with values {1, 2, 3} and variable $X_{2}$ with values {A, B, C}. Suppose the following sample of 10 individuals is observed.
@@
2 3 C
3 1 A
4 1 A
5 3 C
6 3 C
7 1 A
8 3 C
9 1 A
10 3 C
We can see that the probabilities on both histograms $H(X_{1})$ and $H(X_{2})$ are 0.4, 0.1 and 0.5, so that $H(X_{1})=H(X_{2})$. Combinations of values from the two variables in the sample determine a bijection, and we have indiscernible variables with $SU(X_{1},X_{2})=1$ showing a perfect correlation.
@@
Indeed, to see this, just consider two categorical random variables, $X$ and $Y$, with categories $\{x_{1},x_{2}\}$ and $\{y_{1},y_{2}\}$, and probabilities $\{p_{1},1-p_{1}\}$ and $\{q_{1}=p_{1}+\epsilon,q_{2}=1-q_{1}\}$. Then, we see that $H(X|Y)=0$, and thus
$SU(X,Y)=2\frac{H(X)}{H(X)+H(Y)}.$
Since $H(X)$ and $H(Y)$ are continuous functions of $p_{i}$ and $q_{i}$, we can make $H(X)+H(Y)\approx 2H(X)$, resulting in 1SU(X,Y)1-SU(X,Y) as small as we want by letting $\epsilon\to 0$. ∎
@@ 4.1 Formal definitions
Indeed, let $A^{\prime}\colon D\to\Sigma_{A^{\prime}}$ and $B^{\prime}\colon D\to\Sigma_{B^{\prime}}$ be two categorical random variables equivalent to $A$ and $B$, respectively. Then, the bijections $\Sigma_{A^{\prime}}\to\Sigma_{A}$ and $\Sigma_{B^{\prime}}\to\Sigma_{B}$ yield a bijection $\Sigma_{A^{\prime}}\times\Sigma_{B^{\prime}}\to\Sigma_{A}\times\Sigma_{B}$. Also, since $A\sim A^{\prime}$, we have $P_{A}=P_{A^{\prime}}$ and similarly, $P_{B}=P_{B^{\prime}}$. It follows that the joint probability measure $P_{A^{\prime}*B^{\prime}}$ and $P_{A*B}$ are also equal. Thus, we get that $A^{\prime}*B^{\prime}$ and ABA*B are equivalent. ∎
@@
Consider $A],[B],[C]\in\overline{\cal C}$, and let us pick the representatives
$A\colon D\to\Sigma_{A};\quad B\colon D\to\Sigma_{B};\quad C\colon D\to\Sigma_{C}.$
Then, we have $B*C:D\to\Sigma_{B}\times\Sigma_{C}$, as well as $A*B\colon D\to\Sigma_{A}\times\Sigma_{B}$. Then, we get
@@
and also
$(A*B)*C\colon D\to(\Sigma_{A}\times\Sigma_{B})\times\Sigma_{C}=\Sigma_{A}\times\Sigma_{B}\times\Sigma_{C}.$
Finally, we also have
$\begin{split}[A*(B*C)](p)&=(A(p),(B*C)(p))=(A(p),(B(p),C(p)))\\ &=((A(p),B(p)),C(p))=[(A*B)*C](p).\end{split}$ (20)
Hence, we get A(BC)=(AB)CA*(B*C)=(A*B)*C. Passing to equivalence classes, we get the associativity
[A]([BC])=([A][B])[C].[A]*([B*C])=([A]*[B])*[C].
For the commutativity, let us observe that $B*A\colon D\to\Sigma_{B}\times\Sigma_{A}$. Hence, we have a bijection between $\Sigma_{A*B}$ and $\Sigma_{B*A}$. Also, for the probability mass function we have
$P_{A*B}(x,y)=P(A=x\wedge B=y)=P(B=y\wedge A=x)=P_{B*A}(y,x).$
Thus, we get that ABA*B and BAB*A are equivalent. Hence
[A][B]=[AB]=[BA]=[B][A],[A]*[B]=[A*B]=[B*A]=[B]*[A],
as claimed. ∎
@@
Now we observe that $\Sigma_{A*\Phi}=\Sigma_{A}\times\{\phi\}$, so we have a bijection $\Sigma_{A}\to\Sigma_{A*\Phi}$ given by $x\mapsto(x,\phi)$. Also, the probability mass function of $\Phi$ is $P_{\Phi}(\phi)=1$, thus we get
$P_{A*\Phi}(x,\phi)=P(A=x|\Phi=\phi)\cdot P(\Phi=\phi)=P(A=x)=P_{A}(x).$
Hence, $A*\Phi$ and $A$ have the same histogram.
Thus, we conclude that $A*\Phi$ is equivalent to $A$. It follows that
$A]*[\Phi]=[A*\Phi]=[A],$
as claimed. ∎

### v2

@@ 1 Introduction
Symmetric Uncertainty has been introduced in [6], and it is defined as
@@
where $X,Y$ are categorical random variables, (i.e.: random variables taking values in a finite set, not necessarily numerical), $H(\cdot)$ is the Shannon entropy [10], and $H(\cdot,\cdot)$ is the joint Shannon entropy.
Generalization of SU for the multivariate case, as well as its properties, and, in particular, its potential to define a distance metric among categorical random variables have been explored in [5, 4, 12, 11]. Indeed, [4] bring us to realize that entropic correlation is an example of similarity.
@@
Our work is organized as follows.
In Section 2, we review some basic concepts from probability and information theory in order to understand the motivation and definition of the symmetric uncertainty.We also discuss the concept of entropy from two perspectives: one related to categorical random variables and another related to partitions of a set. Finally, we conclude this section with some technical lemmas concerning entropy and joint entropy that will be needed later.
In Section 3, we define an appropriate quotient space of categorical random variables and endow it with a metric structure based on entropy and symmetric uncertainty.
In Section 4, we endow the same quotient space of categorical random variables with a monoid structure and establish the compatibility of the topological and algebraic structures.
Finally, in Section 5, we offer some concluding remarks highlighting the potential usefulness of our formalization for statistical practitioners.
@@ 2 Background: Entropic Correlation and Symmetric Uncertainty
Let us consider a fixed probability space $(D,\Omega_{D},P)$, and let $X\colon D\to\Sigma_{X}$ be a categorical (non-numeric, nominal or cualitative) random variable with possible values on the finite set $\Sigma_{X}=\{x_{1},\ldots,x_{k}\}$ and probability mass function $P_{X}\colon\Sigma_{X}\to\mathbb{R}$ defined as $P_{X}(x_{i}):=P(X=x_{i}):=P(X^{-1}(x_{i}))$. The entropy $H$ of the variable $X$ is a measure of the uncertainty in predicting the value of $X$, or size of the uncertainty on $X$, and is defined as [9]:
$H(X):=-\sum_{i}P_{X}(x_{i})\log_{2}(P_{X}(x_{i})).$ (2)
Any base could be used for the log, but base 2 is used for historical reasons connected to information theory [10]. $H(X)$ can also be interpreted as a measure of the amount of information a discrete random variable $X$ requires for describing its behavior; or more compactly, as the variety inherent to $X$.
Given another categorical random variable $Y\colon D\to\Sigma_{Y}$, with possible outcomes in the set $\Sigma_{Y}=\{y_{1},y_{2},\ldots,y_{m}\}$, we may consider the conditional probability
$P(x_{i}|y_{j}):=\frac{P(X^{-1}(x_{i})\cap Y^{-1}(y_{j}))}{P(Y^{-1}(y_{j}))}.$
Then, we can define the conditional entropy
$H(X|Y):=-\sum_{j}\left[P_{Y}(y_{j})\sum_{i}P(x_{i}|y_{j})\log_{2}(P(x_{i}|y_{j}))\right].$ (3)
The conditional entropy $H(X|Y)$ quantifies the amount of information needed to describe the outcome of $X$ given that the value of $Y$ is known.
@@
For more details, we refer the reader to [1], page 15.
Mutual Information. Using entropy and conditional entropy introduced above, the mutual information  [^1] In certain literature mutual information is denoted as $I(X;Y)$; see [3]. $MI(X|Y)$ is introduced [7]. Mutual Information measures the reduction in uncertainty about the value of $X$ when the value of $Y$ is known, and is defined as:
@@
and since $H(X,Y)$ is symmetric by direct substitution of $X,Y$ into equation (1), it results that $MI(X|Y)$ is symmetric as well, a quite convenient property for a paired measure. A formal proof can be found, for example, as Theorem 1 in [3].
Symmetric Uncertainty. On the other hand, like with other informational association measures, MI tends to increase its value when the number of categories in $X$ and/or $Y$ increases, that is, it is biased towards high cardinality features [8]. Therefore, seeking to compensate for such bias, MI has to be normalized by the sum of entropies of the features. This measure, called Symmetric Uncertainty [6] is expressed as:
@@
As a final observation regarding $SU$, we recall from [12] the following results.
@@
$H({\cal Q}|{\cal R}):=-\sum_{Q\in{\cal Q}}\sum_{R\in\cal R}P(Q\cap R)\log\left(\frac{P(Q\cap R)}{P(R)}\right).$ (10)
@@
Lemma 2.2 (Lemma 9.1.5 in [14]).
@@ 3.1 Measuring similarity on a dataset: between observations and between variables
A set of data, or dataset, consists of $m$ observations taken across $n$ variables, or put another way, of $n$ random variables $X_{j}$ evaluated on $m$ subjects $q_{i}$. Let us set $x_{ij}=X_{j}(q_{i})$, so we can arrange the observed values as shown in Table 1.
@@
Looking row-wise at the dataset, any two observations (rows) can be compared to determine their similarity. If two rows are equal throughout all their values, we say that they are in full similarity or equivalently, we say that the distance between these observations is zero. If not all the values on the two rows are equal, several techniques are available to compute distances in this $n$-dimensional space, where a mix of data types is not unusual; see for instance Suárez-Díaz et al [13].
We can also look at the same dataset in a column-wise manner [4]. In exploratory experiments and in model construction, one tries to establish how the value in one column (the outcome or class) results from the values in other columns, revealing a concomitance or correlation between the random variables.
@@ 20 R I L H D N
Table 2 presents data example for the personality traits of students applying for an internship. Computing probabilities from the observed frrequency, Table 3 shows the calculation of $SU$ for the variables Neatness and Hired based on the three terms in Equation (8).
Number of cases = 20
Neatness PP(Neatness) log PP(Neatness)
U 0.25 -2
S 0.25 -2
R 0.5 -1
HH(Neatness) = 1.5
Hired PP(Hired) log PP(Hired)
N 0.55 -0.8625
Y 0.45 -1.1520
HH(Hired) = 0.9928
Neatness, Hired PP(Neatness, Hired) log PP(Neatness, Hired)
U, N 0.1 -3.3219
U, Y 0.15 -2.7370
S, N 0.1 -3.3219
S, Y 0.15 -2.7370
R, N 0.35 -1.5146
R, Y 0.15 -2.7370
HH(Neatness, Hired) = 2.4261
SUSU(Neatness, Hired) = 0.0535
Table 3: Calculating SUSU(Neatness, Hired) from Table 2.
The full set of resulting $SU$ values between each of the variables and the Hired response are:
  • SUSU(Neatness, Hired) = 0.0535
  • SUSU(Creativity, Hired) = 0.4627 (largest computed $SU$)
  • SUSU(Punctuality, Hired) = 0.0535
  • SUSU(IQuotient, Hired) = 0.0535
  • SUSU(AttentionType, Hired) = 0.0192
@@ 3.2 Entropic equivalence and the metric structure
Less standardized is the concept of a similarity metric. Here we recall the definition given by Chen et al in [2].
@@
In [2], the authors also gave conditions for a similarity metric to yield a distance metric. More concretely, they established a precise relationship between normalized distances and similarity metrics. Here, normalized means that both the distance and the similarity map into the interval $0,1]\subset\mathbb{R}$. For the reader convenience, let us recall their result here.
Lemma 3.2 (Corollary 1 in [2]).
@@
Let $\cal C$ be a set of categorical random variables defined on a sample space $D$. If $X\colon D\to\Sigma_{X}$ and $Y\colon D\to\Sigma_{Y}$ are two elements in $\cal C$}, we say that they are indiscernible if there is a bijection $h\colon\Sigma_{X}\to\Sigma_{Y}$ and $Y=h\circ X$ almost everywhere.
Let us notice that, for two indiscernible categorical random variables $X$ and $Y$, their corresponding partition $\cal X$} and $\cal Y$} are equal almost everywhere. We will denote this by ${\cal X}\sim{\cal Y}$. In particular, the relative entropy given by Equation  (10) is zero for two indiscernible categorical random variables.
The indiscernibility of categorical random variables introduces an equivalence relation in $\cal C$. Let us denote by $\overline{{\cal C}}$ the space of equivalence classes. Notice that any two equivalence classes may have the same number of categories or a different number of categories.
@@
Example of indiscernibles. Consider variable $X$ with values {1, 2, 3} and variable $Y$ with values {A, B, C}. Table 5 shows a sample of 10 individuals. Considering the probability given by the observed frequency, we can see that the histogram of $X$ is 0.5, 0.1, 0.4, while that of $Y$ is 0.4, 0.1, 0.5. Here, we can define the bijection $h\colon\{1,2,3\}\to\{A,B,C\}$ given by $h(1)=C;h(2)=B;h(3)=A$, and then the partition defined on the domain by $Y$ and $h\circ X$ are the same.
@@
2 1 C
3 3 A
4 3 A
5 1 C
6 1 C
7 3 A
8 1 C
9 3 A
10 1 C
@@
Let $X$ be a categorical random variable, and let us consider $Y$ as a noisy copy of $X$, namely: $Y=X$ with probability $1-\epsilon$, and an independent redraw of the categories with probability $\epsilon$. Thus, $Y$ is almost a function of $X$, which means that knowing $X$ almost determines $Y$. Then, as $\epsilon\to 0$ we have $H(Y|X)\to 0$ and $H(Y)\to H(X)$, hence
$SU(X,Y)=2\frac{H(Y)-H(Y|X)}{H(X)+H(Y)}\to 2\frac{H(X)}{H(X)+H(X)}=1.$
It follows that $d(X,Y)=1-SU(X,Y)\to 0.$
@@ 4.1 Formal definitions
Notice that, in the context of the previous definition, for each $(a,b)\in\Sigma_{A}\times\Sigma_{B}$ we have
$C^{-1}(a,b)=A^{-1}(a)\cap B^{-1}(b),$
[…]
${\cal C}={\cal A}\cap{\cal B}.$
@@
Indeed, let $A^{\prime}\colon D\to\Sigma_{A^{\prime}}$ and $B^{\prime}\colon D\to\Sigma_{B^{\prime}}$ be two categorical random variables equivalent to $A$ and $B$, respectively. Then, the bijections
$h_{A}\colon\Sigma_{A^{\prime}}\to\Sigma_{A}\quad\text{and}\quad h_{B}\colon\Sigma_{B^{\prime}}\to\Sigma_{B}$
[…]
@@
Consider $A],[B],[C]\in\overline{\cal C}$. For any representatives
$A\colon D\to\Sigma_{A};\quad B\colon D\to\Sigma_{B};\quad C\colon D\to\Sigma_{C},$
we need to show that
@@
and
$(A*B)*C\colon D\to(\Sigma_{A}\times\Sigma_{B})\times\Sigma_{C}=\Sigma_{A}\times\Sigma_{B}\times\Sigma_{C},$
are indiscernible.
We clearly have the bijection between the codomains, and also for the corresponding partition we have
${\cal A}\cap({\cal B}\cap{\cal C})=({\cal A}\cap{\cal B})\cap{\cal C}.$
For the commutativity, let us observe that $A*B\colon D\to\Sigma_{A*B}=\Sigma_{A}\times\Sigma_{B}$ and $B*A\colon D\to\Sigma_{B*A}=\Sigma_{B}\times\Sigma_{A}$. Hence, we have a bijection between $\Sigma_{A*B}$ and $\Sigma_{B*A}$, and for the partitions we clearly have
${\cal A}\cap{\cal B}={\cal B}\cap{\cal A}.$
Thus, we have [A][B]=[B][A][A]*[B]=[B]*[A]. ∎
@@
Now we observe that $\Sigma_{A*\Phi}=\Sigma_{A}\times\{\phi\}$, so we have a bijection $\Sigma_{A}\to\Sigma_{A*\Phi}$ given by $x\mapsto(x,\phi)$.
Also, the partition corresponding to $\Phi$ is the trivial one ${\cal D}=\{D\}$, and thus the partition of $A*\Phi$ is
${\cal A}\cap{\cal D}={\cal A}.$
Hence, we have $A]*[\Phi]=[A],$ as claimed. ∎